1)Two planets with masses M and 4M are at a distance R each other. A meteor with mass m entered to their gravitational field. At what distance from the larger planet on the line joining the planets will the net gravitational force on the mass m be zero


1
Expert's answer
2021-05-09T13:04:49-0400
"G\frac{mM}{(R-x)^2}=G\frac{4mM}{x^2}\\x=\frac{2}{3}R"